the measure of an angle is 25 more than its supplement. what is the measure of the larger angle?

Answer by Theo(13342)

x equals the angle.
180 - x equals the supplement.
x = (180 - x) + 25
simplify to get x = 205 - x
add x to both sides of the equation to get 2x = 205
solve for x to get x = 102.5
the supplement is 180 - 102.5 = 77.5
their sum is 102.5 + 77.5 = 180
their difference is 102.5 - 77.5 = 25
solution is confirmed to be good.
solution is that the larger angle is equal to 102.5 degrees.